OpenStudy (anonymous):

The y-intercept of the line whose equation is x - y = -1 is

OpenStudy (anonymous):

is it 1

myininaya (myininaya):

To find y-intercept, set x equal 0 and solve for y.

myininaya (myininaya):

Yep.

OpenStudy (anonymous):

help me with this The line through (1, 0) and parallel to x - y = 7.

myininaya (myininaya):

You can find the slope of a line by putting the equation into y=mx+b form where m is the slope. Can you arrange your equation x-y=7 into y=mx+b form? The reason we want to do this is because lines that are parallel have the same slope.

OpenStudy (anonymous):

still confused

myininaya (myininaya):

Can you write x-y=7 into y=mx+b form? Basically I'm asking you to solve x-y=7 for y.

OpenStudy (***[isuru]***):

 so the slope of both graphs r equal.. which mean the line parallel to x - y = 7 is somewhat something which looks like x - y = c ( c = intercept) now substitute the given coordinates (1 , 0) x = 1 y = 0 1 - 0 = c c =1 so the function is.. x - y = 1 y = x -1 got it ?